# MADation: Face Morphing Attack Detection with Foundation Models

### v2

@@ 3.3 MAD Baselines
Text-Image (TI) MAD: FMs like CLIP have demonstrated exceptional zero-shot learning performance across several downstream tasks, including action recognition in videos, sentiment analysis and car model classification [45]. Taking this into account, we evaluate the selected FM zero-shot learning performance on MAD, by simultaneously using its text and image encoders. To this end, TI processes image-text pairs where the textual input specifies the two possible classification labels. The predicted label is determined by analyzing the similarity score between the image embedding and the input text embeddings. However, the potential of this technique should not be overestimated based on its success in general downstream domains such as the ones specified above given the domain-specific nature of MAD. The fact that no adaption is performed to the specific requirements of the downstream MAD task makes TI prone to underperform when compared with methods that consider MAD’s specific characteristics, such as MADation.
@@
Feature Extractor (FE) MAD: To determine the importance of the network adaption allowed by LoRA in MADation, we develop an experiment where the FM is not fine-tuned. In this scenario, the FM works as a frozen feature extractor on top of which a fully-connected layer is trained to perform classification, using the binary cross-entropy loss as in Equation 5. This experiment allows us to assess the suitability of the FM’s original feature space to discriminate between MAD classes while quantitatively measuring the improvements introduced by adapting the FM’s weights to the domain-specific downstream MAD task with LoRA.
@@ 4 Experimental Setup
Datasets: The Synthetic Morphing Attack Detection Development (SMDD) dataset [14] was selected as the training dataset of the proposed models. SMDD [14] is a synthetic-based MAD dataset, consisting of 25k bona-fide images generated using the StyleGAN2-ADA framework [33, 52] and 15k morphing attacks created from the bona-fide samples using the OpenCV morphing technique [37]. The SMDD dataset was chosen for training as it ensures privacy by avoiding the use of real face images and because it has shown remarkable success in MAD solution development [13, 38, 7] and public competitions [27]. The benchmarking datasets proposed in these works, MAD22 [27] and its extension MorDIFF [13], were also chosen to ensure the results’ comparability and data domain and identity disjoint from the SMDD training data. The evaluation benchmarks are based on the Face Research Lab London (FRLL) dataset [19] and thus contain the same 204 bona-fide images. Additionally, the same image pairs were used to create the morphing attacks in MAD22 and MorDIFF. The MAD22 dataset includes morphed images from five different approaches: three image-level techniques (FaceMorpher, OpenCV [37], and Webmorph) and two GAN-based representation-level methods (MIPGAN I and II [56]). The morphing samples of the MorDIFF dataset were generated with a diffusion autoencoder [42]. In addition to MAD22 [27], the FRLL-Morphs dataset [48] is utilized for evaluation so that a comparison is possible with methods that are not evaluated on MAD22. The FRLL-Morphs dataset [48], similarly derived from the Face Research London Lab dataset [19], serves as a benchmark for MAD evaluation. The dataset comprises 204 genuine samples and over 1,000 morphed faces per technique, generated using five distinct morphing methods: Style-GAN2 [33, 52], WebMorph [18], AMSL [39], FaceMorpher [43], and OpenCV [37].
Image Pre-Processing: Before being used as an input to the FM, each sample is cropped following [14] and then resized to $224\times 224$ pixels to comply with the image resolution originally used to train CLIP [45]. During training, all samples are also subject to data augmentation using random horizontal flipping, following [50, 38]. Due to the success achieved by tokenization in foundation models’ NLP applications, Alexey et al. [2] proposed preprocessing images into tokens before feeding them to the FM’s image encoder. All the experiments mentioned in this document follow this tokenization process, represented in Figure 1. Each input sample is divided into non-overleaping regions, fed to a linear projection layer [2]. The resultant patch embeddings are appended to a learnable embedding, the class (CLS) token [2], which constitutes an image representation that helps classify the input into predefined categories. Additional position embeddings are considered to preserve the spatial order of the original sample’s patches. The final embedding vector is then fed to the image encoder [2].
@@
Implementation Details: All models were trained for 40 epochs using AdamW optimizer [35] with momentum of 0.9 and weight decay of 0.05 [11]. For MADation, we used ViT-B and ViT-L architectures with model and header learning rates of 1e-5 and 1e-4, respectively. The LoRA parameters were set to $r=2$, $\alpha=\{4,8\}$ and $dropout=\{0.4,0.2\}$ for ViT-B and ViT-L respectively. For FE, we used a header learning rate of 1e-2 for both architectures. ViT-FS models were trained with learning rates of 1e-5/5e-5 (model/header) for ViT-B and 1e-4/1e-4 for ViT-L. The batch size was set to 256 for all settings.
Evaluation Metrics: The MAD evaluation metrics were selected to ensure compliance with the ISO/IEC 30107-3 [28] standard and enable consistent benchmarking and comparability with previous studies [27, 13]. Performance is reported using the Bona-fide Presentation Classification Error Rate (BPCER), the Attack Presentation Classification Error Rate (APCER), and the detection Equal Error Rate (EER). The BPCER quantifies the proportion of bona-fide images misclassified as attack samples, while the APCER measures the proportion of attack images misclassified as bona-fide samples. The detection EER is the error rate at the operating point where the BPCER and APCER are equal, offering a concise measure of the system’s overall performance balance. To cover different operational points and present comparative results, we report both the APCER at fixed BPCER values and the BPCER at fixed APCER values, evaluated at values of 1.0%, 10%, and 20%.
@@ 5 Results and Discussion
[…]
@@
[…]
@@
[…]
@@
Baselines toward MADation (ViT-FS and FE): We further explore two alternative approaches, ViT-FS and FE. ViT-FS is trained from scratch, making it possible to assess the potential of the underlying ViT architectures for the MAD task without relying on the built-in knowledge of FMs. FE makes use of this knowledge by using CLIP as a frozen feature extractor on top of which a binary fully-connected layer is trained to perform classification, allowing us to assess the suitability of the FM’s original feature space to discriminate between MAD classes. The results achieved by both approaches are presented in Table 1. When the ViT-B architecture is considered, FE outperforms ViT-FS in 4 of the 6 considered benchmarking datasets, with an average EER difference of 0.68 pp.. For ViT-L, FE largely surpasses ViT-FS performance, improving the average EER by 8.83 pp.. Similar tendencies can be observed for metrics such as APCER@BPCER=20% and APCER@BPCER=10%. The superiority of FE when compared to ViT-FS possibly derived from the large number of trainable parameters of ViT-B and ViT-L. These large-scale networks require large amounts of training data to properly learn the considered task without overfitting, which might undermine ViT-FS capacity to learn the MAD task given the reduced size of the SMDD dataset. On the other hand, FE benefits from the FM’s previous knowledge, which was acquired during a pre-training phase with a massive amount of training data, which justifies its superior performance. Furthermore, the fact that this performance difference is higher when using ViT-L (8.83 pp. vs 0.68 pp. difference on average EER when using ViT-L and ViT-B, respectively) mostly derives from ViT-FS decreased performance in this scenario, which further reinforces that the limited capacity of ViT-FS to learn the MAD task is strongly correlated with the large number of trainable parameters of the considered networks. Hence, it can be concluded that taking advantage of the built-in knowledge of the FM results in better performance than training the network from scratch with a reduced amount of training data. Nonetheless, FE’s MAD performance still has the potential for further enhancement, making it worth exploring whether adapting the FM to the downstream MAD task results in increased performance.
@@
MADation: As previously discussed, FMs can generalize to a wide variety of downstream tasks but can show limited capacity when handling domain-specific tasks such as MAD. The previously analyzed FE results highlight this characteristic, since the EER values still leave space for improvement, revealing that the FM’s feature space is most likely significantly misaligned for the MAD task. Although it is also possible that the selected classification network has saturated its capacity given the FM’s feature space and should thus be deeper, we argue that the problem is most likely arising from the feature space misalignment which can be corrected through FM’s adaption, as highlighted in previous studies [54, 8, 11]. Furthermore, this type of adaption allows the network to have more flexibility without giving up on the knowledge acquired during the FM’s pre-training phase, which might constitute a good trade-off between the properties of FE and ViT-FS. Hence, we propose to adapt CLIP to the MAD task with LoRA layers, resulting in our proposed approach, MADation. We evaluate MADation using the same benchmarks as FE and ViT-FS, to allow for a fair comparison with these approaches, as shown in Table 1. It can be seen that MADation achieves the best and/or second-best performance levels for most of the evaluated metrics and benchmarks. In particular, ViT-B is the best-performing method on average in 2 out of the 7 evaluated metrics, and the second-best-performing method on the remaining 5. For these 5 metrics, ViT-L presents the best overall performance. The analysis of the average EER also reveals MADation’s superiority when compared with the remaining approaches, as it surpasses ViT-FS and FE by 10.24 pp. and 9.56 pp., respectively, for ViT-B and by 16.93 pp. and 8.10 pp., respectively, for ViT-L. MADation’s improvements when compared with FE prove the importance of performing a correct FM adaption to downstream domain-specific tasks such as MAD. Simultaneously, MADation’s superiority regarding ViT-FS shows that the network adaption provided by LoRA does not prevent the final network from benefitting from the knowledge acquired during pre-training. Hence, it is possible to conclude that MADation reaches an efficient trade-off between preserving the FM’s built-in knowledge and fine-tuning it to the downstream task, resulting in improved MAD performance.
@@
Comparison with the recent MAD approaches: To further extend our study and verify if MADation shows competitive performance with recent MAD approaches, we compare our proposed framework with several MAD architectures previously proposed in the literature. The comparison considered all the reported results in the literature that complied with the SYN-MAD 2022 competition [27] by training on SMDD [14] and testing on the MAD22 [27] (and its derivatives [13]) as well as the works trained on SMDD and tested on the FRLL-Morphs [48] evaluation benchmark. This might have missed comparisons to some published MAD techniques (that did not follow this protocol or are not publicly available) but provides a wide comparison with many of them and relies on a public and clear benchmark. In Table 2, we start by comparing the MAD techniques [14, 13, 46, 30, 27, 44] with public results on MAD22 [27] and its extension MorDIFF [13], including some of the submitted solutions to the SYN-MAD 2022 competition [27]. Our approach presents the best and/or second-best performance in 33 out of the 42 evaluated scenarios. Note that no other approach in Table 2 evaluated BPCER at a fixed APCER for MorDIFF. In particular, ViT-B presents remarkable results in FaceMorph and OpenCV, while ViT-L consistently outperforms the remaining techniques in FaceMorph, MIPGAN II and WebMorph. Although the D-FW [44] approach presents very competitive performance levels for all available metrics, it should be kept in mind that these models use a multi-task learning framework that incorporates 3D facial information. This leads to a significant increase in computational demands and justifies the achieved performance levels. Furthermore, this information could also be included in the proposed MADation approach and would likely result in increased performance. Nonetheless, applications such as these fall out of the scope of the current work, which focuses on providing a computationally inexpensive solution to the MAD task through the usage of LoRA layers to adapt a pre-trained FM. We further extend our study to the well-known FRLL-Morphs dataset [48] in Table 3, to present a more comprehensive comparison with previous works [38, 7, 14, 3]. It can be seen that MADation scores first and/or second place in 16 out of the 20 evaluated scenarios, with ViT-B achieving the best performance levels in 10 of these scenarios. In particular, ViT-B achieves the lowest EER for FRLL-WebMorph and FRLL-AMSL, reducing the previously best EER values by 0.59 pp. and 0.15 pp., respectively. Overall, MADation presents competitive performances with a wide set of recent MAD solutions, highlighting the importance of exploring FM’s potential in biometrics tasks such as MAD.
@@ 6 Conclusion
Acknowledgments: This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.

### v3

@@ 3.3 MAD Baselines
Text-Image (TI) MAD: FMs like CLIP have demonstrated exceptional zero-shot learning performance across several downstream tasks, including action recognition in videos, sentiment analysis and car model classification [45]. Taking this into account, we evaluate the selected FM zero-shot learning performance on MAD, by simultaneously using its text and image encoders. To this end, TI processes image-text pairs where the textual input specifies the two possible classification labels. The predicted label is determined by analysing the similarity score between the image embedding and the input text embeddings. However, the potential of this technique should not be overestimated based on its success in general downstream domains such as the ones specified above given the domain-specific nature of MAD. The fact that no adaption is performed to the specific requirements of the downstream MAD task makes TI prone to underperform when compared with methods that consider MAD’s specific characteristics, such as MADation.
@@
Feature Extractor (FE) MAD: To determine the importance of the network adaption allowed by LoRA in MADation, we develop an experiment where the FM is not fine-tuned. In this scenario, the FM works as a frozen feature extractor on top of which a fully connected layer is trained to perform classification, using the binary cross-entropy loss (Equation 5). This experiment allows us to assess the suitability of the FM’s original feature space to discriminate between MAD classes while quantitatively measuring the improvements introduced by adapting the FM’s weights to the domain-specific downstream MAD task with LoRA.
@@ 4 Experimental Setup
Datasets: The Synthetic Morphing Attack Detection Development (SMDD) dataset [14] was selected as the training dataset of the proposed models. SMDD [14] is a synthetic-based MAD dataset, consisting of 25k bona-fide images generated using the StyleGAN2-ADA framework [33, 52] and 15k morphing attacks created from the bona-fide samples using the OpenCV morphing technique [37]. The SMDD dataset was chosen for training as it ensures privacy by avoiding the use of real face images and because it has shown remarkable success in MAD solution development [13, 38, 7] and public competitions [27]. The benchmarking datasets proposed in these works, MAD22 [27] and its extension MorDIFF [13], were also chosen to ensure the results’ comparability and a data domain identity disjoint from the SMDD training data. The evaluation benchmarks are based on the Face Research Lab London (FRLL) dataset [19] and thus contain the same 204 bona-fide images. Additionally, the same image pairs were used to create the morphing attacks in MAD22 and MorDIFF. The MAD22 dataset includes morphed images from five different approaches: three image-level techniques (FaceMorpher, OpenCV [37], and Webmorph) and two GAN-based representation-level methods (MIPGAN I and II [56]). The morphing samples of the MorDIFF dataset were generated with a diffusion autoencoder [42]. In addition to MAD22 [27], the FRLL-Morphs dataset [48] is utilized for evaluation so that a comparison is possible with methods that are not evaluated on MAD22. The FRLL-Morphs dataset [48], similarly derived from the Face Research London Lab dataset [19], serves as a benchmark for MAD evaluation. The dataset comprises 204 genuine samples and over 1,000 morphed faces per technique, generated using five distinct morphing methods: Style-GAN2 [33, 52], WebMorph [18], AMSL [39], FaceMorpher [43], and OpenCV [37].
Image Pre-Processing: Before being used as an input to the FM, each sample is cropped following [14] and then resized to $224\times 224$ pixels to comply with the image resolution originally used to train CLIP [45]. During training, all samples are also subject to data augmentation using random horizontal flipping, following [50, 38]. Due to the success achieved by tokenization in FMs’ NLP applications, Alexey et al. [2] proposed preprocessing images into tokens before feeding them to the FM’s image encoder. All the experiments mentioned in this document follow this tokenization process, represented in Figure 1. Each input sample is divided into non-overleaping regions which are fed to a linear projection layer [2]. The resultant patch embeddings are appended to a learnable embedding, the class (CLS) token [2], which constitutes an image representation that helps classify the input into predefined categories. Additional position embeddings are considered to preserve the spatial order of the original sample’s patches. The final embedding vector is then fed to the image encoder [2].
@@
Implementation Details: All models were trained for 40 epochs using the AdamW optimizer [35] with momentum of 0.9 and weight decay of 0.05 [11]. For MADation, we used ViT-B and ViT-L architectures with model and header learning rates of 1e-5 and 1e-4, respectively. The LoRA parameters were set to $r=2$, $\alpha=\{4,8\}$ and $dropout=\{0.4,0.2\}$ for ViT-B and ViT-L respectively. For FE, we used a header learning rate of 1e-2 for both architectures. ViT-FS models were trained with learning rates of 1e-5/5e-5 (model/header) for ViT-B and 1e-4/1e-4 for ViT-L. The batch size was set to 256 for all settings.
Evaluation Metrics: The MAD evaluation metrics were selected to ensure compliance with the ISO/IEC 30107-3 [28] standard and enable consistent benchmarking and comparability with previous studies [27, 13]. Performance is reported using the Bona-fide Presentation Classification Error Rate (BPCER), the Attack Presentation Classification Error Rate (APCER), and the detection Equal Error Rate (EER). The BPCER quantifies the proportion of bona-fide images misclassified as attack samples, while the APCER measures the proportion of attack images misclassified as bona-fide samples. The detection EER is the error rate at the operating point where the BPCER and APCER are equal, offering a concise measure of the system’s overall performance balance. To cover different operational points and present comparative results, we report both the APCER at fixed BPCER values and the BPCER at fixed APCER values, evaluated at values of 1%, 10%, and 20%.
@@ 5 Results and Discussion
Average 11.89 missing39.36 missing17.18 missing12.39 42.64 missing22.14 missing13.97
Worst 33.37 missing82.97 missing55.18 43.92 94.12 72.55 52.94
@@
Average 20.04 74.13 40.56 23.06 missing60.16 30.57 20.54
Worst missing26.40 88.84 56.80 missing37.80 71.57 missing41.67 missing29.90
MADation (ours) FaceMorph 0.40 0.40 0.00 0.00 0.49 0.00 0.00
MIPGAN_I 20.32 55.88 29.08 20.32 79.41 35.78 15.69
@@
WebMorph 20.40 47.60 20.40 20.40 82.35 37.25 13.24
MorDIFF 19.26 48.40 24.45 19.26 84.47 34.95 15.53
Average missing11.94 29.24 14.05 11.04 60.40 18.90 7.57
Worst 20.40 55.88 29.08 missing20.40 100.00 37.25 15.69
@@
Baselines toward MADation (ViT-FS and FE): We further explore two alternative approaches, ViT-FS and FE. ViT-FS is trained from scratch, making it possible to assess the potential of the underlying ViT architectures for the MAD task without relying on the built-in knowledge of FMs. FE makes use of this knowledge by using CLIP as a frozen feature extractor on top of which a binary fully connected layer is trained to perform classification, allowing us to assess the suitability of the FM’s original feature space to discriminate between MAD classes. The results achieved by both approaches are presented in Table 1. When the ViT-B architecture is considered, FE outperforms ViT-FS in 4 of the 6 considered benchmarking datasets, with an average EER difference of 0.68 pp.. For ViT-L, FE largely surpasses ViT-FS performance, improving the average EER by 8.83 pp.. Similar tendencies can be observed for metrics such as APCER@BPCER=20% and APCER@BPCER=10%. The superiority of FE when compared to ViT-FS possibly derives from the large number of trainable parameters of ViT-B and ViT-L. These large-scale networks require large amounts of training data to properly learn the considered task without overfitting, which might undermine ViT-FS’ capacity to learn the MAD task given the reduced size of the SMDD dataset. On the other hand, FE benefits from the FM’s previous knowledge, which was acquired during a pre-training phase with a massive amount of training data, which justifies its superior performance. Furthermore, the fact that this performance difference is higher when using ViT-L (8.83 pp. vs 0.68 pp. difference on average EER when using ViT-L and ViT-B, respectively) mostly derives from ViT-FS decreased performance in this scenario, which further reinforces that the limited capacity of ViT-FS to learn the MAD task is strongly correlated with the large number of trainable parameters of the considered networks. Hence, it can be concluded that taking advantage of the built-in knowledge of the FM results in better performance than training the network from scratch with a reduced amount of training data. Nonetheless, FE’s MAD performance still has the potential for further enhancement, making it worth exploring whether adapting the FM to the downstream MAD task results in increased performance.
@@
MADation: As previously discussed, FMs can generalize to a wide variety of downstream tasks but can show limited capacity when handling domain-specific tasks such as MAD. The previously analysed FE results highlight this characteristic, since the EER values still leave space for improvement, revealing that the FM’s feature space is most likely significantly misaligned for the MAD task. Although it is also possible that the selected classification network has saturated its capacity given the FM’s feature space and should thus be deeper, we argue that the problem is most likely arising from the feature space misalignment which can be corrected through FM’s adaption, as highlighted in previous studies [54, 8, 11]. Furthermore, this type of adaption allows the network to have more flexibility without giving up on the knowledge acquired during the FM’s pre-training phase, which might constitute a good trade-off between the properties of FE and ViT-FS. Hence, we propose to adapt CLIP to the MAD task with LoRA layers, resulting in our proposed approach, MADation. We evaluate MADation using the same benchmarks as FE and ViT-FS, to allow for a fair comparison with these approaches, as shown in Table 1. It can be seen that MADation achieves the best and/or second-best performance levels for most of the evaluated metrics and benchmarks. In particular, ViT-B is the best-performing method on average in 2 out of the 7 evaluated metrics, and the second-best-performing method on the remaining 5. For these 5 metrics, ViT-L presents the best overall performance. The analysis of the average EER also reveals MADation’s superiority when compared with the remaining approaches, as it surpasses ViT-FS and FE by 10.24 pp. and 9.56 pp., respectively, for ViT-B and by 16.93 pp. and 8.10 pp., respectively, for ViT-L. MADation’s improvements when compared with FE prove the importance of performing a correct FM adaption to downstream domain-specific tasks such as MAD. Simultaneously, MADation’s superiority regarding ViT-FS shows that the network adaption provided by LoRA does not prevent the final network from benefitting from the knowledge acquired during pre-training. Hence, it is possible to conclude that MADation reaches an efficient trade-off between preserving the FM’s built-in knowledge and fine-tuning it to the downstream task, resulting in improved MAD performance.
@@
Comparison with the recent MAD approaches: To further extend our study and verify if MADation shows competitive performance with recent MAD approaches, we compare our proposed framework with several MAD architectures previously proposed in the literature. The comparison considered all the reported results in the literature that complied with the SYN-MAD 2022 competition [27] by training on SMDD [14] and testing on the MAD22 [27] (and its derivatives [13]) as well as the works trained on SMDD and tested on the FRLL-Morphs [48] evaluation benchmark. This might have missed comparisons to some published MAD techniques (that did not follow this protocol or are not publicly available) but provides a wide comparison with many of them and relies on a public and clear benchmark. In Table 2, we start by comparing the MAD techniques [14, 13, 46, 30, 27, 44] with public results on MAD22 [27] and its extension MorDIFF [13], including some of the submitted solutions to the SYN-MAD 2022 competition [27]. MADation presents the best and/or second-best performance in 23/21 (ViT-B/ViT-L) out of the 42 evaluated scenarios. Note that no other approach in Table 2 evaluated BPCER at a fixed APCER for MorDIFF. In particular, ViT-B presents remarkable results in FaceMorph and OpenCV, while ViT-L consistently outperforms the remaining techniques in FaceMorph, MIPGAN II and WebMorph. Although the D-FW [44] approach presents very competitive performance levels for all available metrics, it should be kept in mind that these models use a multi-task learning framework that incorporates 3D facial information. This leads to a significant increase in computational demands and justifies the achieved performance levels. Furthermore, this information could also be included in the proposed MADation approach and would likely result in increased performance. Nonetheless, applications such as these fall out of the scope of the current work, which focuses on providing a computationally inexpensive solution to the MAD task through the usage of LoRA layers to adapt a pre-trained FM. We further extend our study to the well-known FRLL-Morphs dataset [48] in Table 3, to present a more comprehensive comparison with previous works [38, 7, 14, 3]. It can be seen that MADation scores first and/or second place in 12/9 (ViT-B/ViT-L) out of the 20 evaluated scenarios. In particular, ViT-B achieves the lowest EER for FRLL-WebMorph and FRLL-AMSL, reducing the previously best EER values by 0.59 pp. and 0.15 pp., respectively. Overall, MADation presents competitive performances with a wide set of recent MAD solutions, highlighting the importance of exploring FM’s potential in biometrics tasks such as MAD.
@@ 6 Conclusion
Acknowledgments: This research work has been funded by the German Federal Ministry of Education and Research, and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE, and the Slovenian ARIS research programs Metrology and Biometric Systems (P2-0250) and Computer Vision (P2-0214).